I need help in figuring out the answer to these two problems. I am having a tough time trying to figure them out. Can someone please help me with this. Thank you in advance for your help.

1) Use the distributive law to multiply. 4(a+b)

a) 8ab

b) 4a+b

c) 4ab

d) 4a+4b




2) Use the distributive law to factor the given expression. 6a+3b

a) 6(a+b)

b) 2(3a+3b)

c) 3(2a+b)

d) 3(2a+3b)

Answer by user_dude2008(1862)

Dist Law: a(b+c)=ab+ac


4(a+b) = 4a+4b ----> answer is d



6a+3b = 3(2a+b) ----> answer is c